(12 , —4) and slope : -3

Found 2 solutions by ikleyn, ewatrrr:
Answer by ikleyn(52788)

.

    y - (-4) = -3*(x-12).


Or any other EQUIVALENT equation.

Solved.

---------------------

It is very elementary and basic skill that every student must develop to the level to solve such problems on his  (or her)
own  AUTOMATICALLY  without asking assistance from outside.

    Similar to as you answer  "4"   when you are asked  "How many is  2*2  ?".

To help you to develop such skills,  the special lesson was created in this site

    - Equation for a straight line having a given slope and passing through a given point

Find there a complete explanation on how to do it and learn the subject once and for all.



Answer by ewatrrr(24785)


Hi
(12 , —4) and slope : -3
 y + 4 = -3(x - 12)
   y = -3x + 32  (in slope-intercept form)
Wish You the Best in your Studies.